OpenStudy (anonymous):

Snookers Lumber can convert logs into either lumber or plywood. In a given day, the mill turns out three times as many units of plywood s lumber. It makes a profit of $20 on a unit of lumber and $50 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of $17680

OpenStudy (anonymous):

17680 = 20x + 50y y = 3x where y = plywood units x = lumber units so 17680 = 20x + 50(3x) 17680 = 20x + 150x 17680 = 170x 104 = x Lumber units = 104 Plywood units = 312

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

y=3x (y=plywood, x= lumber) 20x+50y=17680 20x+50(3x)=17680 20x+150x=17680 170x=17680 x=104 y= 3(104) x=104, y=312

OpenStudy (anonymous):

thanks